OpenStudy (anonymous):

Hello I need help please (x+5) (x-6)

OpenStudy (anonymous):

x^2-x-30

OpenStudy (ivettef365):

FOIL

OpenStudy (raden):

use the foil (x+5) (x-6) = x(x-6) + 5(x-6) = x^2 - 6x + 5x - 30 simplify similar terms = x^2 - x - 30

OpenStudy (anonymous):

what is foil?

OpenStudy (anonymous):

Thank you guys I got it now thank you :o)